OpenStudy (anonymous):

A square painting is surrounded by a frame. The outside edges of the frame are x inches in length and there is a 4-inch border between the painting and the frame. What is the total area of the border?

OpenStudy (anonymous):

so, could you guys solve this?

OpenStudy (faiqraees):

The answer should come in terms of x right?

OpenStudy (anonymous):

–16x – 64 –8x + 16 16x – 64 x2 + 16x + 64

OpenStudy (faiqraees):

Area of border = Total area - Area of Painting Area of border = x^2- (x-8)^2 Now expand it and you've your area. If you want to know how i got (x-8)^2, let me know

OpenStudy (anonymous):

that's not one of the possible answers

OpenStudy (anonymous):

–16x – 64 –8x + 16 16x – 64 x2 + 16x + 64

OpenStudy (faiqraees):

I know I haven't given you the answers, Expand the brackets and then simplify the equation and you will have your answer

OpenStudy (anonymous):

16x - 64?

OpenStudy (anonymous):

Can we see a diagram?

OpenStudy (faiqraees):

Yep correct

OpenStudy (anonymous):

@FaiqRaees Could you explain how you got the answer?

OpenStudy (faiqraees):

Area of border = Total area - Area of Painting Area of border = x^2- (x-8)^2 Area of border = x^2- (x^2 -16x+64) Area of border = x^2- x^2 +16x-64 Area of border = 16x-64